What is the molar concetration that results from a 2.75 M solution of potassium nitrate that is tenfold serially diluted by a factor of 10 taised to the 7th?
1
Expert's answer
2019-05-27T06:09:43-0400

D.F. = C1(KNO3) / C2(KNO3)

C2(KNO3) = C1(KNO3) / D.F. = 2.75 M / 10 = 0.275 M

C3(KNO3) = C2(KNO3) / D.F. = 0.275 M / 7 = 0.04 M